Exercise 14.69 Calculate the pH of each solution. Submit My Answer, Give Up Correct Part C (H30 6.2x10-6 M Express your answer using two decimal places. pH 5.21 Submit My Answers Give Up Correct Part D H O 6.4x10 M Express your answer using two decimal places. pH My Answers Give Up Submit
Answer #1

Part C

[H3O+] = 6.2 * 10-6 M

pH = -log [H3O+]

pH = -log (6.2 * 10-6) = 5.21

Part D

[H3O+] = 6.4 * 10-4 M

pH = -log [H3O+]

pH = -log (6.4 * 10-4) = 3.19